# DP-LLM: Runtime Model Adaptation with Dynamic Layer-wise Precision Assignment

### v1

@@ 5.2 Asynchronous Estimation
Although lightweight, DP-LLM’s relative error estimation lies in the critical path of inference, as it requires the immediate input for estimation. To further minimize slowdown, as depicted in Figure 6, DP-LLM performs asynchronous estimation by using the previous input vector instead of the immediate one whenever possible. Recent studies dejavu ; infinigen have demonstrated that activations in Transformer architectures change slowly across blocks due to residual connections. This creates an opportunity for layers directly connected to residual connections to use the previous residual output as the input for error estimation. Specifically, query, key, value, and up-projection correspond to such layers. The latency of asynchronous estimation can be masked by overlapping it with other layer computations.
@@ 6.2 Inference Latency Evaluation
We measure the inference latency overhead introduced by DP-LLM’s precision selector. Specifically, we implement a PyTorch-based inference pipeline and optimize it using the torch.compile feature torch_compile for latency measurement. To isolate the overhead of the precision selector and exclude the impact of varying effective bitwidth, we fix the effective bitwidth and compare the latency against a static precision allocation baseline (LLM-MQ) without the precision selector. The evaluation is conducted on two hardware platforms: NVIDIA Jetson Orin AGX 64GB jetson and NVIDIA RTX 4060 Ti 16GB rtx .
@@ 6.3 Per-Query QoS Validation
Since DP-LLM aims to match the overall average precision to the target precision on a best-effort basis, individual queries may exhibit deviations from the target precision. To assess the impact of DP-LLM on query-level QoS, we collect the distribution of effective bitwidth when inferencing queries in Alpaca dataset alpaca , which consists of handwritten instructions for chatbot. The results, summarized in Table 7, report the increase of the 90th and 99th percentile effective bitwidth relative to the mean effective bitwidth. Even at the 99th percentile, the geomean increase remains below 3%, demonstrating that DP-LLM has minimal adverse impact on per-query QoS.
@@ 7 Conclusion
We propose DP-LLM, a runtime model adaptation scheme that enables dynamic layer-wise precision assignment. Based on our observation that the layer-wise sensitivity to precision varies across decoding iterations, DP-LLM devises a novel mechanism that assigns precision dynamically at each decoding iteration based on a per-layer relative error threshold and a lightweight precision selector. DP-LLM achieves state-of-the-art performance across a wide range of datasets and configurations by leveraging the unrecognized opportunity of dynamism in sensitivity.
@@ Appendix A Formulation of Layer-wise Maximum Precision Selection
When the weights of each layer $W_{i}$ are quantized to a $b$-bit weight $W_{i,b}$, the objective function to minimize can be expressed as follows. Here $N$ denotes the number of layers, $B$ is the set of available precisions, $M_{i}$ is the memory usage of the $i$th layer, and $b_{targ}$ is the target precision.
@@ B.1 Setup and Methodology
To ensure full GPU saturation during the decoding phase, each model is compiled using PyTorch’s torch.compile function. This compiles a CUDA Graph of the model, eliminating kernel launch overhead and improving GPU utilization efficiency. Latency is measured by generating 100 tokens across 10 repetitions, with the average effective bitwidth set equal to the target precision. This bitwidth is fixed by setting each layer’s threshold to either infinity or negative infinity.
@@ B.3 Fine-tuning Results
Both Llama-3-8B and Phi-3-Medium are fine-tuned on a single RTX 3090 GPU with 24GB of VRAM. Llama-3-8B completes fine-tuning in approximately one hour, utilizing 14GB of VRAM, while Phi-3-Medium takes about two hours with 21GB of VRAM usage. On an A100 80GB GPU, Llama-3-8B requires around 30 minutes, and Phi-3-Medium takes approximately one hour to complete fine-tuning.

### v3

@@ 5.2 Asynchronous Estimation
Although lightweight, it is undesirable that DP-LLM’s relative error estimation, which requires the immediate input of each layer, lies on the critical path of inference. To minimize slowdown, as depicted in Figure 6, DP-LLM performs asynchronous estimation by using the previous input vector instead of the immediate one whenever possible. Recent studies dejavu ; infinigen have demonstrated that activations in Transformer architectures change slowly across blocks due to residual connections. This creates an opportunity for layers directly connected to residual connections to use the previous residual output as the input for error estimation. Specifically, query, key, value, and up-projection correspond to such layers. The latency of asynchronous estimation can be masked by overlapping it with other layer computations.
@@ 6.2 Inference Latency Evaluation
We measure the inference latency overhead introduced by DP-LLM’s precision selector. Specifically, we implement DP-LLM on top of gpt-fast gpt-fast , which optimizes it using the torch.compile feature torch_compile . To isolate the overhead of the precision selector and exclude the impact of varying effective bitwidth, we fix the effective bitwidth and compare the latency against a static precision allocation baseline (LLM-MQ) without the precision selector. The evaluation is conducted on two hardware platforms: NVIDIA Jetson Orin AGX 64GB jetson and NVIDIA RTX 4060 Ti 16GB rtx .
@@ 6.3 Per-Query QoS Validation
Since DP-LLM aims to match the overall average precision to the target precision on a best-effort basis, individual queries may exhibit deviations from the target precision. To assess the impact of DP-LLM on query-level QoS, we collect the distribution of effective bitwidth when inferencing queries in the Alpaca dataset alpaca , which consists of handwritten instructions for chatbot. The results, summarized in Table 7, report the increase of the 90th and 99th percentile effective bitwidth relative to the mean effective bitwidth. Even at the 99th percentile, the geomean increase remains below 3%, demonstrating that DP-LLM has minimal adverse impact on per-query QoS.
@@ 7 Conclusion
We propose DP-LLM, a runtime model adaptation scheme that enables dynamic layer-wise precision assignment. Based on our observation that the layer-wise sensitivity to precision varies across decoding steps, DP-LLM devises a novel mechanism that assigns precision dynamically at each decoding step based on a per-layer relative error threshold and a lightweight precision selector. DP-LLM achieves state-of-the-art performance across a wide range of datasets and configurations by leveraging the unrecognized opportunity of dynamism in sensitivity.
Acknowledgments
This work was supported by Samsung Electronics Co., Ltd. and also by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (No. 2021-0-02068, Artificial Intelligence Innovation Hub) and the National Research Foundation of Korea (NRF) grant (RS-2024-00405857), both funded by the Korea government (MSIT).
@@ Appendix A Formulation of Layer-wise Maximum Precision Selection
When the weights of each layer $W_{i}$ are quantized to a $b$-bit weight $W_{i,b}$, the objective function to minimize can be expressed as Equation 6. Here, $N$ denotes the number of layers, $B$ is the set of available precisions, $M_{i}$ is the memory usage of the $i$th layer, and $b_{targ}$ is the target precision.
@@ B.1 Setup and Methodology
To ensure full GPU saturation during the decoding phase, by building on top of gpt-fast [38], each model is compiled using PyTorch’s torch.compile function. This compiles a CUDA Graph of the model, eliminating the kernel launch overheads and improving GPU utilization efficiency. Latency is measured by generating 100 tokens across 10 repetitions, with the average effective bitwidth set equal to the target precision. This bitwidth is fixed by setting each layer’s threshold to either infinity or negative infinity.
@@ B.3 Fine-tuning Results
Both Llama-3-8B and Phi-3-Medium are fine-tuned on a single RTX 3090 GPU with 24GB of VRAM. Llama-3-8B completes fine-tuning in approximately one hour, utilizing 14GB of VRAM, while Phi-3-Medium takes about two hours with 21GB of VRAM usage. On an A100 80GB GPU, Llama-3-8B requires about 30 minutes, and Phi-3-Medium takes approximately one hour to complete fine-tuning.